OpenStudy (anonymous):

Probability Question! Please Help! There are 7 yellow, 3 red and 5 green balls in a bag. Jack draws two balls at random, without replacing them again in the bag. What is the probability that both balls that he picks are Yellow?

terenzreignz (terenzreignz):

What's the probability that the FIRST ball he picks is yellow? :)

OpenStudy (anonymous):

Now what I want to ask is that, see when he first draws out a ball, that probability that he'll pick a Yellow ball is 7/15, correct, now he takes his hand inside the bag and again draws out a ball, so the probabilty of the next event that he'll draw out a yellow will be 7/14? correct? Because he has already picked up one ball? so the total balls that remain in the bag is 14? But the probabilty of picking up a Yellow doesnt change, correct?

OpenStudy (anonymous):

Because we dont know whether he picked out a yellow or any other color in the first draw...

terenzreignz (terenzreignz):

Ok, you're right in saying that the odds of the first ball being yellow is 7/15 that is because there are 7 yellow balls out of 15 balls in all... As for the second ball, we assume he picked a yellow ball in the first draw, and then there would only be 6 yellow balls left out of 14... so?

OpenStudy (anonymous):



OpenStudy (anonymous):

Why do we assume that he picked the yellow in the first draw??? W don't know that! May be he picked some other color

terenzreignz (terenzreignz):

Then if you assume he picks another color, then there is no chance that both balls are yellow, so why bother with those possibilities?

OpenStudy (anonymous):

I don't get it ... :(

terenzreignz (terenzreignz):

By saying that the odds of the first ball being yellow is 7/15, the fact that it isn't 1, that already says that we acknowledge that the first ball MIGHT NOT be yellow So we're just going assuming that the first ball is already yellow... :) Do you get it now? :)

OpenStudy (anonymous):

It's difficult to digest.

terenzreignz (terenzreignz):

Yeah, statistics isn't my cup of tea neither :P Sorry I couldn't explain it any better :(

OpenStudy (anonymous):

That's okay. :(

OpenStudy (apoorvk):

Okay I am an idiot, what was I thinking all this time? okay this is simple: probability of drawing the first ball yellow = 7/15 now 14 balls left, and 6 yellow ball among those, so probability of picking another yellow ball = 6/14 hence probability = 7/15 x 6/14